OpenStudy (anonymous):

is this correct? q1) ((lim)/(x→∞)) ((7+4p²)/(11p²-19)) = ((p²((7/(p²))+4))/(p²(11-((19)/(p²))))) =(4/(11)) q2) ((lim)/(t→1)) ((4t²-4)/(t²-t)) =((8t)/(2t-1)) =((8(1))/(2(1)-1))=(8/1)=8

hartnn (hartnn):

yes, both are correct

hartnn (hartnn):

4 (t+1)(t-1)/t(t-1) = 4(t+1)=4*2=8 i would have solved 2nd like that^ .

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^